# Improving sampling of binding free energy differences between covalently bound ligands in alternate binding pockets using MT-REXEE

### v1

@@ Abstract
The primary limitation for the application of alchemical free energy methods to a wider variety of complex molecular systems in achieving reasonable sampling. Flexible binding complexes often have high free energy barriers, which require prohibitively long simulations to sample sufficiently to obtain reliable free energy estimates. An example of such a system is the FabB–acyl chain complex, which features acyl chain substrates from 4 to 16 carbons in length. Previous experimental evidence supports that there are also at least two accessible binding pockets for the acyl chain bound to FabB. Sampling of both pockets would require prohibitively long MD simulations or the creation of complicated and nonobvious collective variables. With the multiple topology replica exchange of expanded ensemble (MT-REXEE) enhanced sampling approach, we can obtain highly efficient sampling between both pockets by adaptively growing and shrinking the chains in the simulation ensemble, allowing each simulation to visit chain lengths where transitions between the pockets do occur.
To enable this enhanced sampling for the chain growth problem, we demonstrate a new and improved conformational swapping selection algorithm for MT-REXEE, and validate this algorithm on a previously studied simultaneous calculation of multiple relative binding free energies. Using the new swapping approach gives significantly enhanced sampling even for this simpler problem, as demonstrated by faster convergence of free energy estimates.
This case study demonstrates the utility of MT-REXEE and its open-source implementation for systems that feature high free energy barriers for a subset of ligands of interest, demonstrating a valuable addition to the existing stable of enhanced sampling methods.
@@ 1 Introduction
Alchemical free energy (FE) calculations have become standard in many computational molecular researchers’ workflows, but their application is primarily limited to systems with comparatively little conformational flexibility 1, 2. Though the primary application of FE simulations has historically been small-molecule drug discovery 3, 4, their utility extends far beyond such binding free energy calculations. However, these less standard applications often involve systems that exhibit significant conformational flexibility, either in the protein, the ligand, or both, requiring prohibitively long simulations to converge FE estimates. In many cases, systems with significant flexibility can remain trapped in metastable states for longer than is reasonable to simulate them. These limitations have prompted the discovery of numerous enhanced sampling methods, including our own multiple topology replica exchange of expanded ensemble 5.
Protein and enzyme engineering is an important, emerging application for FE calculations in large and flexible systems 6, 7. Often, engineered protein systems involve multimeric complexes, which introduce flexible inter-protein interfaces. One such example is the enzymes within the fatty acid synthesis (FAS) cycle. This cycle is of interest for the industrial biosynthesis of a wide range of products, including bio-fuel precursors 8, 9. FAS involves enzymes that form either dimers or tetramers. The substrate for FAS enzymes is transported between each enzyme in the cycle by the acyl carrier protein (ACP) 10, 11, 12, 13, 14. Thus simulation of the ACP–enzyme complex involves multiple flexible binding interfaces.
The covalent intermediate for FAS is both biologically relevant and a more straightforward state to quantify sampling enhancements than the ACP–enzyme complex. The covalent intermediate forms after the acyl chain substrate has been passed to the CYS residue on the respective FAS enzyme. The covalent intermediate state for FAS enzymes suffers from poor sampling in MD simulations due to the presence of a large, flexible substrate whose conformational flexibility is restricted by the covalent bonding to the protein and a relatively narrow binding pocket.
@@
These same sampling challenges also occur when applying FE simulations to covalent drug development 15, 4, 16, 17 as covalent drugs have a similar covalent intermediate. Covalent drug development is an emerging field given the increased potency and lifetime within the system. However, enhancing the sampling of AFE methods by capturing multiple binding and unbinding events as alchemical parameters change to minimize the configurational free energy barriers is not possible for covalent ligands using molecular mechanics simulations, as covalent bonds cannot be formed or broken. FE calculations with covalent ligands, both in the enzyme engineering and covalent drug cases, thus require alternate methods of achieving enhanced sampling to improve the accuracy and utility of computational methods.
@@
Existing enhanced sampling methods are insufficient for the flexible covalent ligands in the FAS system. Enhanced sampling methods that purely use the alchemical variable present in free energy calculations, $\lambda$, to enhance sampling include Hamiltonian replica exchange 18 and expanded ensemble 19, 20. These methods to enhance sampling require that either the end states or the intermediate states have significantly different free energy barriers in configuration space. For the FAS examples proposed earlier, we are interested in the free energy of binding as a function of substrate chain length, which can vary by 2 carbon intervals from 4 to 16 carbons. Performing a single alchemical transformation from 4 to 16 carbons would not give information about the intermediate chain lengths, and performing iterative transformations would not result in sufficiently low FE barriers between alternate bound configurations at long chain lengths to converge sampling.
@@
Collective variable (CV) dependent methods, including metadynamics 21, 22, umbrella sampling 23, 24, and weighted ensemble approaches 25, represent another approach to enhanced sampling. However, they require significant prior knowledge about the conformational changes that enable transition between trapped metastable states. For the FAS examples, although we know that longer chains must be able to move between pockets, the number of degrees of freedom involved makes it challenging to identify relevant CVs that can describe such a transition involving all of the atoms. Intermediate motions along the proposed CV likely sample high-energy configurations, which decrease sampling efficiency. Even with an approximate path, these CV-dependent methods require prior knowledge of where any alternate pockets may be or significant trials to derive a more complete CV.
@@
We previously presented the enhanced sampling method, multiple topology replica exchange of expanded ensemble (MT-REXEE), which specifically targets enhancing sampling in flexible receptor-ligand binding systems 5. In this work, we present improvements made to the MT-REXEE method through the introduction of a novel swapping algorithm, random range swapping, which increases the efficiency of the calculations and improves FE estimate convergence. This method works by enabling trajectory frames other than the last one in the iteration to be used as potential swappable states rather than utilizing redundant end-states. We also apply MT-REXEE to the difficult sampling problem, calculating the acyl chain length dependence on the binding FE for the FabB-acyl complex. This complex demonstrates the ability of MT-REXEE to significantly enhance conformational sampling in bound systems by coupling free energy simulations together configurationally.
The FabB-acyl complex represents an excellent test case demonstrating how MT-REXEE can accelerate sampling without prior knowledge of the behavior of the system. Previous experimental evidence revealed that the I108M mutation in FabF significantly reduces synthesis of acyl chain products longer than 8 carbons in length 26. The analogous mutation G107M in FabB, however, did not produce the same reduction of activity on medium chain substrates 26. We hypothesize that an alternate binding pocket is accessible in FabB, which enables FabB to continue synthesis even in the event that the canonical binding pocket is blocked by the introduced MET residue. Here we demonstrate the ability of MT-REXEE to sample both the canonical binding pocket (pocket A) and the alternate binding pocket (pocket B) in both WT and G107M FabB variants, without biasing the simulation towards the alternate binding pocket or even identifying it beforehand. With MT-REXEE, not only do we access both pockets within the same simulation, but we can quantify the FE penalty of the G107M mutation in pocket A and the relative stability between of pockets A and B.
@@
Sampling of both distinct binding pockets is not possible using standard MD or independent expanded ensemble simulations within reasonable simulation lengths (<<500 ns) for substrates longer than 4 carbons. But we can leverage the increased conformational flexibility of the short-chain substrates through these conformational swaps to increase sampling in long-chain simulations. We present this method as a suitable method for overcoming high configurational free energy barriers which are present for only a subset of ligands being evaluated. We also demonstrate how MT-REXEE simulations can produce both configurational and alchemical free energy values for all evaluated end points and between two or more distinct binding pockets.
@@ 2 Theory: Separating Alchemical and Conformational Free Energies
One application of the MT-REXEE method is to reconstruct the binding FE dependence on both ligand identity and bound conformation for systems that feature two or more distinct binding locations. We apply the MT-REXEE method to the FabB acyl covalent intermediate complex, as this system features a wide variety of ligand substrates (acyl chain substrates vary from 4 to 16 carbons) and is expected to sample two distinct acyl chain binding locations (A and B). For this system, each MT-REXEE simulation performs one alchemical transformation and can feature two or more configurational states (Figure 1A). MT-REXEE enables conformational exchange between each transformation, enabling more comprehensive sampling of the two or more conformational states. From the set of MT-REXEE simulations, we obtain the dependence of binding FE on both the substrate and the binding conformation (Figure 1B).
Figure 1: (A) A set of MT-REXEE simulations which perform transformations of the acyl substrate from C2 to C4, C4 to C6, and C6 to C8. Each simulation, visiting a range of alchemical states $\lambda$ in [0,1], has configurations that should (in the limit of sufficient sampling) visit both pocket A (orange) and pocket B (brown). The dashed boxes indicate states in each individual simulated transformation, and the dark blue arrows show where configurational swaps can occur, at the alchemical end states. (B) When we separate the transformations by end state and configurational state, we can estimate each of the FE differences we compute from a MT-REXEE simulation. We compute the alchemical FE difference for each transformation independently in each pocket (light blue) as well as the configurational FE difference at at least one end state (pink). This enables us to compute the relative FE difference between any combination of alchemical end state and configurational state.
@@
$\Delta G_{ij}=-k_{B}T\ln{\frac{Q_{j}}{Q_{i}}}=-k_{B}T\ln{\frac{\int_{\Gamma_{j}}e^{-\frac{U_{j}(\vec{q})}{k_{B}T}}d\vec{q}}{\int_{\Gamma_{i}}e^{-\frac{U_{i}(\vec{q})}{k_{B}T}}d\vec{q}}}$ (1)
@@
$\Delta G_{ij,A}=-k_{B}T\ln{\frac{Q_{j,A}}{Q_{i,A}}}=-k_{B}T\ln{\frac{\int_{\Gamma_{j,A}}e^{-\frac{U_{j}(\vec{q})}{k_{B}T}}d\vec{q}}{\int_{\Gamma_{i,A}}e^{-\frac{U_{i}(\vec{q})}{k_{B}T}}d\vec{q}}}$ (2)
@@
There may be concern that the dimensionality of the conformational space defining A and B will vary between transformations as each transformation is performed with different numbers of dummy particles. At this point, it is important to note that this method is being used purely for differences in free energies. This means that an equivalent transformation, between states $i$ and $j$, is performed in solvent as well as in complex with the protein. Since the calculation of $\Delta\Delta G$ is independent for each transformation and we maintain only three bonded interactions between the dummy and ”real” atoms during the transformation, Boresch shows that the dummy atom contributions cancel, assuming equal dummy atom treatment in both solvent and protein complex 27. We could rectify the issue of dimensionality differences in partition space through the addition of dummy atoms to all transformations, such that all atoms necessary for all transformations are present, but many remain in the dummy state for both end-points of the FE transformation. Boresch shows that the configurational free energy (Z) is given by equation 3.
$Z$ $=Z_{J}Z_{vib}Z_{PMF}$ (3)
$Z$ $=Z_{J,D}Z_{J}Z_{vib,D}Z_{vib}Z_{PMF,D}Z_{PMF}$ (4)
In this equation, $Z_{J}$ is the contribution from the Jacobian, $Z_{vib}$ is the contribution from the vibrational motions, and $Z_{PMF}$ is the contribution from the potential of mean force. In the same way that we separate the contribution for the dummy atoms that change alchemical state during the transformation (Equation 4), we can use the same logic to add a second class of dummy atoms that remain in the decoupled alchemical state. Their contribution would be statistically identical for both sides of the free energy cycle and would thus cancel in the same manner.
In this application, the distinction between pocket A and B exists solely in the protein complex conformation and thus in solvent $\Delta G_{ij,A,solv}=\Delta G_{ij,B,solv}$ but this is not necessarily the case so the two will be left as independent variables. Thus the relative binding free energy between alchemical states $i$ and $j$ ($\Delta\Delta G_{ij,A}$) can be given by equation 5 (see Figure 7 for the full thermodynamic cycle). For simplicity, when dealing with multiple subsequent transformations, we instead use the notation $\Delta\Delta G_{n,A}$ where n denotes the transformations in order from the base transformation.
@@
Finally, we need to compute a configurational FE difference between pockets A and B, i.e., $\Delta G_{conf,0}$ (pink arrow) in Figure 1B. In standard MD simulations, which sample from the “native” Boltzmann distribution with no conformational biases, the relative probability of sampling within the conformational space of pocket A versus pocket B is given by the integrated Boltzmann factor over the phase spaces volumes corresponding to the configurations; the configurations where to the chain occupies pocket A versus the set of configurations where the chain occupies in pocket B. for Boltzmann sampling, then this becomes simply the ratio of time (i.e. number of frames) spent in each configurational state. This is shown in Equation 6 where $p_{A}$ and $p_{B}$ and the probability of sampling within the conformational space of pocket A and pocket B respectively, $\beta=k_{B}T$, $k_{B}$ is the Boltzmann factor, $T$ is the temperature, and $\Gamma_{A}$ and $\Gamma_{B}$ are the sets of configurations corresponding to the chain being in pocket A and B respectively.
@@
In this application, we use our EE alchemical free-energy simulations to compute the configurational free energy rather than standard MD, which means we cannot extract the configurational free energy from the entire simulation trajectory to attempt to compute probabilities with the simulation or pocket A and pocket B, as it features non-physical intermediate states. Instead, we can extract a subset of the sampled configurations from the physically meaningful end states ($\lambda=0$), and we can apply equation 6 to determine the configurational free energy difference. We require only one configurational FE difference between states. This choice should be the state for which the kinetics of exchange between configurational states is the fastest. In this application, this is the C2 acyl chain end state. This state had statistically equivalent pocket sampling to that of the independent MD simulations (see Figure 3B), and the uncertainty in this step is less than the uncertainty in the alchemical process 5. Thus we can use the relative probability of the C2 chain sampling either pocket A or B with equation 7 to compute the configuration FE difference $\Delta G_{conf,0}$
This configurational FE difference ($\Delta G_{0,conf}$) along with the alchemical free energy differences in each transformation ($\Delta\Delta G_{n,A}$ and $\Delta\Delta G_{n,B}$) can then be used to compute the FE difference between any end state in any configurational state. Thus, the configurational free energy difference for an end point after $x$ transformations is given by 8 (Figure 15). This relationship will change if the choice of flexible conformational state used to compute the configurational FE difference is different, but we can use any sum of paths which connect the two relevant end states.

### v2

@@ Abstract
The primary limitation for the application of alchemical free energy methods to a wider variety of complex molecular systems is achieving reasonable sampling. Flexible binding complexes often have high free energy barriers, which require prohibitively long simulations to sample sufficiently to obtain reliable free energy estimates. An example of such a system is the complex formed between FabB, an elongating $\beta$-ketoacyl-acyl carrier protein (ACP) synthase (KS) from Escherichia coli and ACP, which carries acyl chains of varying lengths. Previous experimental evidence suggests that growing acyl chains can bind to at least two pockets. With the multiple topology replica exchange of expanded ensemble (MT-REXEE) enhanced sampling approach, we can obtain highly efficient sampling of both pockets by adaptively growing and shrinking the chains in the simulation ensemble, allowing each simulation to visit chain lengths where transitions between the pockets do occur. This enables unbiased sampling of alternate configurational states for large complex systems. Using the new swapping approach gives significantly enhanced sampling even for this simpler problem, as demonstrated by faster convergence of free energy estimates. This case study demonstrates the utility of MT-REXEE and its open-source implementation for systems that feature high free energy barriers for a subset of ligands of interest, demonstrating a valuable addition to the existing stable of enhanced sampling methods.
@@ 1 Introduction
Alchemical free energy (FE) calculations have become standard in many computational molecular researchers’ workflows, but their application is primarily limited to systems with low conformational flexibility 43, 10. Though the primary application of FE simulations has historically been small-molecule drug discovery 1, 30, their utility extends far beyond such binding free energy calculations. These less standard applications, however, often involve systems that exhibit significant conformational flexibility, either in the protein, the ligand, or both, requiring prohibitively long simulations to converge FE estimates. In many cases, systems with significant flexibility can remain trapped in metastable states for longer than is reasonable to simulate them. These limitations have prompted the development of numerous enhanced sampling methods, including our own multiple topology replica exchange of expanded ensemble 14.
Protein and enzyme engineering is an important, emerging application for FE calculations in large and flexible systems 22, 31. Engineered protein systems often involve multimeric complexes, which include flexible inter-protein interfaces. Fatty acid synthases (FASs) provide abundant examples. These biocatalytic assembly lines carry out the iterative biosynthesis of the fatty acids required for enzyme assembly and energy storage, and have served as powerful platforms for the microbial synthesis of oleochemicals  38, 32. Many FAS enzymes form obligate dimers or tetramers and interact with their substrates via acyl carrier proteins (ACPs) 8, 9, 16, 26, 41. Thus, the simulation of the enzyme–ACP complexes involves multiple flexible binding interfaces.
$\beta$-ketoacyl-acyl carrier protein synthases (KSs), which catalyze the carbon-carbon bond formation at the heart of fatty acid synthesis, form a covalent intermediate and provide a nice model system for testing FE methods. Elongation begins when an acyl-ACP associates with a KS, and its acyl cargo is transferred from the PPant arm to a conserved active site cysteine, producing a covalent acyl-KS adduct and holo-ACP 29. Sampling enhancements are more straightforward to quantify for the transacylation state than for the enzyme-ACP complex, yet a change in the stability of this intermediate state could affect the catalytic activity of the enzyme in a similar manner to perturbing the stability of the ACP complex. The absence of an ACP binding interface, as well as the covalent attachment of the substrate in the intermediate, simplifies the analysis for the sake of comparing enhanced sampling methods, as we do not need to consider the delivery of the acyl chain to the enzyme, there is no possibility of the chain leaving, and we do not need to consider the catalytic competency of the protein-protein complex. In the covalent intermediate, sampling issues arise primarily from the presence of a large, flexible substrate whose conformational flexibility is restricted by the covalent bonding to the protein and a narrow primary binding pocket. In a separate study we have performed simulations of the enzyme-ACP complex and present those results in Jiang et al. with a focus on the biophysical interpretation in the context of experimental findings rather than as a verification of the method 20.
@@
These same sampling challenges also occur when applying FE simulations to covalent drug development 12, 30, 36, 37 as covalent drugs also form covalent intermediates with proteins. Covalent drug development is an emerging field, given the potential for increased potency and duration of action 11. However, enhancing sampling of AFE methods by capturing multiple binding and unbinding events as alchemical parameters change to minimize the configurational free energy barriers is not possible for covalent ligands using molecular mechanics simulations, as covalent bonds cannot be formed or broken. FE calculations with covalent ligands, both for enzyme engineering and covalent drug development, thus require alternate methods of achieving enhanced sampling to improve the accuracy and utility of computational methods.
@@
Common enhanced sampling methods have significant disadvantages when applied for the flexible covalent ligands of KSs. One such set of methods uses an alchemical variable present in free energy calculations, $\lambda$, to enhance sampling; these include Hamiltonian replica exchange 19 and expanded ensemble 45, 2 approaches. These methods require that either the end states or the intermediate states have significantly different free energy barriers in configuration space than the purely physical states. However, if we are alchemically modifying an entire fatty acid chain, virtually all of the intermediates will have slow kinetics between alternate configurations because of the many intermediately interacting particles present. In addition, for KS enzymes, we are interested in the free energy of binding as a function of acyl chain length. Performing a single alchemical transformation from short to long chains (e.g. 4 to 16 carbons) would not give information about the intermediate chain lengths, and performing isolated simulations of smaller transformations (i.e. 4 to 6, 6 to 8, etc.) would not result in sufficiently low FE barriers between alternate bound configurations at long chain lengths to converge sampling.
@@
Collective variable (CV) dependent methods, including metadynamics 28, 18, umbrella sampling 44, 40, and weighted ensemble approaches 46, represent alternative approaches to enhanced sampling. However, they require significant prior knowledge about the conformational changes that enable transition between trapped metastable states. For the FAS examples, although we know that longer chains must be able to move when bound, the number of highly correlated degrees of freedom involved makes it challenging to identify relevant CVs that can describe such a transition involving all of the atoms. Intermediate motions along the proposed CV likely sample high-energy configurations, which decrease sampling efficiency. Even with an approximate path, these CV-dependent methods require prior knowledge of where any alternate binding pockets may be or significant trials and sophisticated, problem-specific data analysis to derive a more complete CV. In particular, choices of CV that enhance one type of configurational sampling can eliminate the ability of the system to sample other configurational states not well-described by the given CV.
@@
We previously presented the enhanced sampling method, multiple topology replica exchange of expanded ensemble (MT-REXEE), which specifically targets enhancing sampling in flexible receptor-ligand binding systems 14. In this work, we present improvements made to the MT-REXEE method through the introduction of a novel swapping algorithm, random range swapping, which increases the efficiency of the calculations and improves FE estimate convergence. This method works by enabling trajectory frames other than the final frame in the iteration to be used as potential swappable states in place of the previous solution of simulating redundant end-states. We use this improved algorithm for MT-REXEE to the difficult problem of calculating the binding free energy acyl chain length dependence for the KS-acyl complex. Using this problem, we demonstrate the ability of MT-REXEE to significantly enhance conformational sampling in bound systems by coupling free energy simulations together configurationally.
The elongating KSs from Escherichia coli (E. coli, FabB and FabF 35, provide an excellent test case for demonstrating how MT-REXEE can accelerate sampling. These structurally similar enzymes elongate acyl-ACPs of 4-16 carbons in length. Previous experimental evidence has shown that these enzymes exhibit different responses to analogous mutations. In FabF, the I108M mutation in the acyl binding pocket (pocket A) prevents elongation of acyl chains longer than eight carbons, while in FabB, G107M has a less dramatic effect; it reduces, but still allows, binding of medium chains 24. Crystallographic data suggest that FabB has an alternative binding pocket (pocket B) that allows medium chain binding even when the canonical binding pocket is obstructed by G107M 20. A third pocket (pocket C), which describes acyl chain binding parallel to the Ppant arm, is mainly accessed in the enzyme-ACP complex and appears to serve as a transient binding site for chain delivery to pockets A or B. In this paper, we demonstrate the ability of MT-REXEE to sample pockets A and B in both WT and G107M FabB variants, without biasing the simulation towards the alternate binding pocket. This lack of bias enables the discovery of additional conformational states like pocket C in this instance which may not be prominently sampled, but acts as a proof of concept that MT-REXEE can enable the discovery of novel conformational states. With MT-REXEE, not only do we access both pockets within the same simulation, but we can also quantify the FE penalty of the G107M mutation in pocket A and the relative stability between pockets A and B.
@@
Sampling of both distinct binding pockets is not possible using standard MD or independent expanded ensemble simulations within reasonable simulation lengths (<<500 ns) for substrates longer than 4 carbons. Nonetheless, we can leverage the increased conformational flexibility of the short-chain substrates through these conformational swaps to increase sampling in long-chain simulations. We present this method as a suitable approach when only a subset of ligands have a high configurational free energy barrier to overcome. We also demonstrate how MT-REXEE simulations can produce both configurational and alchemical free energy values for all evaluated end points and between two or more distinct binding pockets.
@@ 2 Theory: Separating Alchemical and Conformational Free Energies
One application of the MT-REXEE method is to reconstruct the binding FE dependence on both ligand identity and bound conformation for systems with two or more distinct binding locations. Here, we apply MT-REXEE to the covalent intermediate of FabB, as this system features a wide variety of ligand substrates (acyl chains of 4-16 carbons) and is expected from crystallographic evidence to sample two distinct acyl chain binding locations (A and B). It is important to note that the expressions below apply for computing the free energy differences between two states regardless of how many possible conformational states there are for the complex (e.g., we ignore the transient binding pocket C). For this system, each MT-REXEE simulation performs one alchemical transformation and can access two or more configurational states (Figure 1A). MT-REXEE enables conformational exchange between each transformation, enabling more comprehensive sampling of the two or more conformational states. From the set of MT-REXEE simulations, we obtain the dependence of binding FE on both the substrate and the binding conformation (Figure 1B).
Figure 1: (A) A set of MT-REXEE simulations that perform transformations of the acyl substrate from C2 to C4, C4 to C6, and C6 to C8. Each simulation, which visits a range of alchemical states $\lambda$ in [0,1], has configurations that should (in the limit of sufficient sampling) visit both pocket A (orange) and pocket B (brown). The dashed boxes indicate states in each individual simulated transformation, and the dark blue arrows show where configurational swaps can occur, at the alchemical end states. (B) When we separate the transformations by end state and configurational state, we can estimate each of the FE differences we compute from a MT-REXEE simulation. We compute the alchemical FE difference for each transformation independently in each pocket (light blue) as well as the configurational FE difference at at least one end state (pink). This enables us to compute the relative FE difference between any combination of alchemical end state and configurational state.
@@
$\Delta G_{ij}=-\frac{1}{\beta}\ln{\frac{Q_{j}}{Q_{i}}}=-\frac{1}{\beta}\ln{\frac{\int_{\Gamma_{j}}e^{-\beta U_{j}(\vec{q})}d\vec{q}}{\int_{\Gamma_{i}}e^{-\beta U_{i}(\vec{q})}d\vec{q}}}$ (1)
@@
$\Delta G_{ij,A}=-\frac{1}{\beta}\ln{\frac{Q_{j,A}}{Q_{i,A}}}=-\frac{1}{\beta}\ln{\frac{\int_{\Gamma_{j,A}}e^{-\beta U_{j,A}(\vec{q})}d\vec{q}}{\int_{\Gamma_{i,A}}e^{-\beta U_{i,A}(\vec{q})}d\vec{q}}}$ (2)
@@
There may be concern that the dimensionality of the conformational space defining A and B will vary between transformations, as each transformation is performed with different numbers of dummy particles. At this point, it is important to note that this method is being used purely for differences in free energies between two environments. In these calculations an equivalent transformation, between states $i$ and $j$, is performed in solvent in addition to the transformation in complex with the protein. Since the calculation of $\Delta\Delta G$ is independent for each transformation and we maintain only three bonded interactions between the dummy and ”real” atoms during the transformation, Boresch shows that the dummy atom contributions cancel, assuming equal dummy atom treatment in both solvent and protein complex 6. We could rectify the issue of dimensionality differences in partition space through the addition of dummy atoms to all transformations, such that all atoms necessary for all transformations are present, but many remain in the dummy state for both end-points of the FE transformation.
In this current application, the distinction between pocket A and B exists solely in the protein complex conformation and thus in solvent $\Delta G_{ij,A,solv}=\Delta G_{ij,B,solv}$ but this is not necessarily the case in general, so the two will be left as independent variables in the derivation for the purpose of generalization. Thus, the relative binding free energy between alchemical states $i$ and $j$ ($\Delta\Delta G_{ij,A}$) can be given by equation 3 (see Figure S2 for the full thermodynamic cycle). For simplicity, when dealing with multiple subsequent transformations, we instead use the notation $\Delta\Delta G_{n,A}$ where n denotes the transformations in order from the base transformation.
@@
Finally, we need to compute a configurational FE difference between pockets A and B, i.e., $\Delta G_{conf,0}$ (pink arrow) in Figure 1B. In standard MD simulations, which sample from the “native” Boltzmann distribution with no conformational biases, the relative probability of sampling within the conformational space of pocket A versus pocket B is given by the integrated Boltzmann factor over the phase spaces volumes corresponding to the configurations; the configurations where to the chain occupies pocket A versus the set of configurations where the chain occupies in pocket B. Assuming proper Boltzmann sampling in the MD simulations, then this ratio of configurational partition functions becomes simply the ratio of time (i.e. number of frames) spent in each configurational state. This is shown in Equation 4 where $p_{A}$ and $p_{B}$ are the probability of sampling within the conformational space of pocket A and pocket B respectively, $\beta=\frac{1}{k_{B}T}$, $k_{B}$ is the Boltzmann factor, $T$ is the temperature, and $\Gamma_{A}$ and $\Gamma_{B}$ are the sets of configurations corresponding to the chain being in pocket A and B respectively.
@@
In the specific application in this paper, we use our EE alchemical free-energy simulations to compute the configurational free energy $\Delta G_{conf,AB}$ rather than using standard MD. Though it may be possible to incorporate the non-physical alchemical intermediate states into the calculation of the configurational free energy difference, for simplicity we restrict the calculation to physical end states. We extract the subset of the sampled configurations from the physical end states ($\lambda=0$ and $\lambda=1$), and we can apply equation 4 to determine the configurational free energy difference. We require only one configurational FE difference between states for one of the ligand identities, as we can use the thermodynamic cycle and the alchemical differences to find the configurational difference for any other chain length. The choice of state to determine the configurational free energy difference should be the state for which the kinetics of exchange between configurational states is the fastest. In the current study, this is the C2 acyl chain end state. This state had statistically equivalent pocket sampling to that of the independent MD simulations (see Figure 3B), and the uncertainty in this step is less than the uncertainty in the alchemical process 14. Thus we can use the relative probability of the C2 chain sampling either pocket A or B with equation 5 to compute the configuration FE difference $\Delta G_{conf,0}$
This configurational FE difference ($\Delta G_{0,conf}$) along with the alchemical free energy differences in each transformation ($\Delta\Delta G_{n,A}$ and $\Delta\Delta G_{n,B}$) can then be used to compute the FE difference between any end state in any configurational state. Thus, the configurational free energy difference for an end point after $x$ transformations is given by 6 (Figure S10). This relationship will change if the choice of flexible conformational state used to compute the configurational FE difference is different, but we can use any sum of paths which connect the two relevant end states.